# 3DSRBench: A Comprehensive 3D Spatial Reasoning Benchmark

### v1

@@ 4.2 Results on 3D Spatial Reasoning Capabilities
LLaVA-v1.5-7B [30] 36.8 38.5 46.4 27.7 31.8
Cambrian-1-8B [44] 44.1 25.6 57.0 36.5 43.1
LLaVA-NeXT-8B [32] 49.6 50.6 62.7 36.8 43.6
@@
Claude-Sonnet [43] 46.9 49.6 60.0 32.8 41.2
Gemini-1.5-Flash [43] 39.2 39.8 59.9 13.2 33.6
Gemini-1.5-Pro [40] 49.1 50.8 62.9 37.5 41.3
GPT-4o-mini [1] 39.1 42.1 51.8 23.4 34.6
GPT-4o [1] 45.3 49.4 62.3 23.0 40.1
@@
Model 3DSRBench-synthetic-common 3DSRBench-synthetic-uncommon
Overall Height Loc. Orient. Multi. Overall Height Loc. Orient. Multi.
Random 20.9 25.0 25.0 16.8 20.1 20.9 25.0 25.0 16.8 20.1
Random++ 45.8 50.0 50.0 41.7 45.0 45.8 50.0 50.0 41.7 45.0
@@
LLaVA-v1.5-7B [31] 42.0 40.0 50.6 20.8 47.6 38.0 41.0 43.6 17.9 45.2
Cambrian-1-8B [44] 48.1 37.5 56.1 39.6 47.6 39.9 35.0 45.7 29.2 41.9
LLaVA-NeXT-8B [32] 45.5 65.0 57.9 10.4 50.0 36.8 47.5 44.5 7.3 46.0
@@
Qwen-VL-Plus [4] 30.7 35.0 37.8 30.2 20.2 21.0 15.0 25.0 22.9 16.1
Qwen-VL-Max [4] 55.2 62.5 69.5 31.2 52.4 48.6 52.5 59.8 24.0 51.6
Claude-Sonnet [3] 47.4 47.5 58.5 26.0 49.2 39.4 60.0 48.2 16.7 38.7
Gemini-1.5-Flash [40] 44.6 57.5 59.8 13.5 44.4 37.7 42.5 45.7 11.5 46.0
Gemini-1.5-Pro [40] 59.9 65.0 69.5 50.0 53.2 49.5 42.5 52.4 40.6 54.8
GPT-4o-mini [1] 46.5 47.5 53.7 36.5 44.4 40.3 42.5 43.9 33.3 40.3
GPT-4o [1] 51.2 70.0 70.1 17.7 46.0 44.3 60.0 58.5 15.6 42.7
@@
We benchmark a wide range of open-sourced and proprietary LMMs on our 3DSRBench real and analyze their 3D spatial reasoning capabilities on different categories of questions. We consider two baseline results: (i) random: a simple baseline that predicts random answers for all visual questions. (ii) random++: a stronger random baseline that predicts consistent answers given different choice orders of a same visual question in CircularEval. We report the full results in Tab. 2. (iii) human: a human-level performance established by human evaluators that did not participate in the data annotation process. We report the full results in Tab. 2.
@@
We make the following observations: (i) State-of-the-art LMMs have limited 3D spatial reasoning capabilities, as found by low performance achieved by state-of-the-art open-sourced and proprietary LMMs, falling far behind human-level performance. (ii) Scaling laws for LMMs are not effective for 3D spatial reasoning. Results show that despite significant more training data and computation spent on the proprietary LMMs, they do not show clear advantages over open-sourced counterparts, featuring high-quality data with efficient training setups. Standard scaling laws demonstrate diminishing returns for 3D spatial reasoning capabilities and we believe more effective approaches, e.g., 3D-aware data, architecture, and training designs, would be necessary to significantly advance 3D spatial reasoning.
4.3 3D Awareness of Vision Encoders
We further study the 3D awareness of vision encoders by benchmarking variants of LLaVA-v1.5-7B [31] on our 3DSRBench with various vision encoders. Specifically, we consider a range of pretrained visual foundation models, i.e., CLIP [39], MAE [21], DINOv2 [38], and SAM [27], and enhance the baseline model with mixed encoders. This can retain the performance on standard benchmarks with the semantic features and potentially involve useful features for 3D reasoning. Next we consider the spatial vision aggregator (SVA) [44] to further fuse the features from multiple vision encoders.
@@
We quantitatively evaluate these LMM variants on our 3DSRBench and report the results in Tab. 3. Results show that with mixed encoders, DINOv2 can improve the overall 3D spatial reasoning abilities of LMM. However, we notice significant improvements for height questions when adopting MAE and SAM as vision encoder, suggesting that having richer visual features could help localize objects better. With SVA [44], we can further improve the LMM with mixed encoder from 37.2% to 37.8%, demonstrating that fusing the semantic features with 3D-aware features from DINOv2 would benefit subsequent reasoning.
@@ 4.4 Robustness to Uncommon Camera Viewpoints
Figure 5: Failure cases of GPT-4o on our 3DSRBench. We find that GPT-4o cannot perform rigorous 3D spatial reasoning and resort to visual cues for reasoning. The bottom two examples show that GPT-4o cannot perform certain complicated 3D spatial reasoning and would produce a short and succinct guess instead.
We study the robustness of 3D spatial reasoning capabilities w.r.t. common and uncommon viewpoints. We evaluate a variety of open-sourced and proprietary LMMs on our 3DSRBench-synthetic-common and 3DSRBench-synthetic-uncommon splits. As demonstrated by the results in Tab. 4, we can see that all LMMs exhibit significantly degraded performance when generalizing from “common” viewpoints to “uncommon” viewpoints, e.g., a 6.9% drop in accuracy for GPT-4o, a 10.4% drop for Gemini-Pro, and a 8.7% drop for LLaVA-NeXT-8B [32]. We visualize two failure examples of GPT-4o in Fig. 4, showing how GPT-4o is capable of predicting the correct answer when prompting with an image from a common viewpoint but fails when being asked about the same question but the image is rendered from an uncommon viewpoint of the exact same scene.
We attribute such degraded performance on “uncommon” viewpoints to two factors: (i) shift in camera 6D viewpoint distribution between training data and images in our 3DSRBench-synthetic-uncommon, and (ii) state-of-the-art LMMs adopt an implicit representation of 3D scenes. Current LMMs are heavily built on the scaling law of data-driven approaches. By leveraging large-scale data for pre-training, alignment, and instruction tuning, models develop an implicit 3D awareness and enable subsequent 3D spatial reasoning capabilities. Despite their success on a wide range of academic and empirical benchmarks, our results demonstrate that state-of-the-art LMMs, open-sourced or proprietary, face severe challenges generalizing to less represented data, which in our case, are images from uncommon camera 6D viewpoints. Notably, although images from these “uncommon” viewpoints are scarce in current datasets, they are in fact very common for specific downstream tasks, such as autonomous navigation and robotic manipulation, where the cameras are often placed on robots rather than held in hand by humans. Degraded performance of current LMMs would largely limit their potential in these areas.
@@ 4.5 Failure Cases
We present some failure cases of GPT-4o in Fig. 5. We make the following two observations: (1) GPT-4o cannot perform rigorous 3D spatial reasoning and resort to various visual cues for reasoning. For the top figure, GPT-4o fail to estimate a rough distance of the two objects, while relying on empirical facts that objects closer to the camera demonstrate more detailed appearances. For the second figure GPT-4o represent orientations with “facing toward approaching traffic”, which is a vague and ineffective orientation representation, leading to the false predicted answer. (2) GPT-4o cannot perform certain complicated 3D spatial reasoning. When being asked about complicated 3D spatial reasoning questions, e.g., ones that require 3D spatial reasoning over 3D locations and orientations, GPT-4o would give a short and succinct answer without complex reasoning. This can further serve as a criteria to study the decision making of GPT-4o, e.g., when GPT-4o skip reasoning and predict answer directly, and analyze its weakness.
@@ 5 Conclusions
In this work we study the 3D spatial reasoning capabilities of LMMs. We present a new benchmark, 3DSRBench, by manually annotating 2,100 visual question-answer pairs on MS-COCO images, featuring diverse and open-vocabulary entities, such as logos on a car or arrows on a billboard, and a balanced data distribution for robust evaluation. To study the robustness of 3D spatial reasoning capabilities w.r.t. camera 6D viewpoints, we further annotate 672 visual question-answer pairs on synthetic multi-view images, each with a common and an uncommon camera viewpoint. Our benchmark further serves as an important diagnosis benchmark for developing 3D spatially intelligent LMMs, shedding light on downstream tasks that require 3D spatial reasoning. Experimental results on different splits of our 3DSRBench provide valuable findings and insights that will benefit future research on 3D spatially intelligent LMMs.
Supplementary Material
A Baseline LMMs
@@
To analyze the 3D spatial reasoning capabilities of state-of-the-art LMMs, we explore a variety of proprietary LMMs, e.g., Qwen-VL, Claude, Gemini, and GPT. We follow Duan et al. [15] and formulate all questions as multiple choice questions with circular evaluation enabled.
@@
B Data Statistics
As discussed in Sec. 3, our 3DSRBench consists of three splits, a real split with 2,100 visual question-answer pairs manually annotated on diverse and open-vocabulary objects from the COCO dataset [29], and two synthetic splits synthetic-common and synthetic-uncommon with 336 visual question-answer pairs in each split, featuring same questions on a 3D scene rendered from a common and an uncommon viewpoint.
Our 3DSRBench consists of 12 question types from four main categories, as demonstrated in Figure 1. All question types are equally distributed, e.g., 175 questions for each question type in 3DSRBench-real.
To show that our 3DSRBench is balanced in answers, we visualize the distribution of closed-form answers in Fig. 6, i.e., one of the four directions (front, left, etc.) or yes/no answers rather than choosing from two given open-vocabulary entities.
C Example Questions in 3DSRBench
@@
Figure 6: Distribution of closed-form answers in our 3DSRBench.

### v2

@@ 4.2 Results on 3D Spatial Reasoning Abilities
LLaVA-v1.5-7B [38] 38.1 39.1 46.9 28.7 34.7
Cambrian-1-8B [52] 42.2 23.2 53.9 35.9 41.9
LLaVA-NeXT-8B [40] 48.4 50.6 59.9 36.1 43.4
InternVL2.5-8B [16] 50.9 45.9 68.1 38.7 43.3
QWen2.5-VL-7B [7] 48.4 44.1 62.7 40.6 40.5
@@
Claude-3.5V-Sonnet [4] 48.2 53.5 63.1 31.4 41.3
Gemini-2.0-Flash [23] 49.8 49.7 68.9 32.2 41.5
Gemini-2.0-Flash-bbox [23] 47.5 45.2 66.5 27.7 41.4
Gemini-2.0-Flash-think [23] 51.1 53.0 67.1 35.8 43.6
GPT-4o-mini [29] 39.7 44.3 52.4 21.0 36.5
GPT-4o [29] 44.2 53.2 59.6 21.6 39.0
QWenVLMax [7] 52.0 45.1 70.7 37.7 44.8
@@
Model 3DSRBench-synthetic-common 3DSRBench-synthetic-uncommon Rel. Drop
Overall Height Loc. Orient. Multi. Overall Height Loc. Orient. Multi. $\delta$
@@
LLaVA-v1.5-7B [39] 42.0 40.0 50.6 20.8 47.6 38.0 41.0 43.6 17.9 45.2 -9.5%
Cambrian-1-8B [52] 48.1 37.5 56.1 39.6 47.6 39.9 35.0 45.7 29.2 41.9 -17.0%
LLaVA-NeXT-8B [40] 45.5 65.0 57.9 10.4 50.0 36.8 47.5 44.5 7.3 46.0 -19.1%
@@
Qwen-VL-Plus [6] 30.7 35.0 37.8 30.2 20.2 21.0 15.0 25.0 22.9 16.1 -31.6%
Qwen-VL-Max [6] 55.2 62.5 69.5 31.2 52.4 48.6 52.5 59.8 24.0 51.6 -12.0%
Claude-Sonnet [5] 47.4 47.5 58.5 26.0 49.2 39.4 60.0 48.2 16.7 38.7 -16.9%
Gemini-1.5-Flash [51] 44.6 57.5 59.8 13.5 44.4 37.7 42.5 45.7 11.5 46.0 -15.6%
Gemini-1.5-Pro [51] 59.9 65.0 69.5 50.0 53.2 49.5 42.5 52.4 40.6 54.8 -32.2%
GPT-4o-mini [29] 46.5 47.5 53.7 36.5 44.4 40.3 42.5 43.9 33.3 40.3 -13.3%
GPT-4o [29] 51.2 70.0 70.1 17.7 46.0 44.3 60.0 58.5 15.6 42.7 -13.5%
@@
We benchmark a wide range of open-sourced and proprietary LMMs on our 3DSRBench real and analyze the 3D spatial reasoning abilities on different types of questions. We consider three baseline results: (i) random: a simple baseline that predicts random answers for all visual questions. (ii) random++: a stronger random baseline that predicts consistent answers given different choice orders of a same visual question in CircularEval. (iii) human: a human-level performance established by human evaluators that did not participate in the data annotation process. We report the full results in Tab. 2.
@@
We make the following observations: (i) State-of-the-art LMMs have limited 3D spatial reasoning capabilities, as found by low performance achieved by state-of-the-art open-sourced and proprietary LMMs, falling far behind human-level performance. (ii) Scaling laws for LMMs are not effective for 3D spatial reasoning. Results show that despite significant more training data and computation spent on the proprietary LMMs, they demonstrate limited advantages over open-sourced counterparts, featuring high-quality data with efficient training setups. Standard scaling laws demonstrate diminishing returns for 3D spatial reasoning abilities and we believe more effective approaches, e.g., 3D-aware data, architecture, and training, would be necessary to significantly advance 3D spatial reasoning.
Design choices of visual encoder.
@@
We study how design choices of visual encoders can benefit 3D spatial reasoning abilities. Built on LLaVA-v1.5-7B [39], we experiment on a range of models with different choices of visual foundation models, i.e., CLIP [47], MAE [25], DINOv2 [46], SAM [34], or model designs, i.e., mixed encoders and visual projectors. Results in Tab. 3 show that with mixed encoders, DINOv2 can improve the overall 3D spatial reasoning abilities of LMMs, specifically for orientation and multi-object reasoning questions that build heavily on object 3D orientations. We also notice significant improvements for height questions when adopting MAE and SAM as vision encoder, suggesting that having richer visual features could help localize objects better. With spatial vision aggregator (SVA) [52], we can further improve the LMM with mixed encoder from 37.2% to 37.8%, demonstrating that fusing the semantic features with 3D-aware features from DINOv2 would benefit subsequent reasoning.
Scaling of language model size.
We study how the scaling of language model, i.e., in terms of the number of parameters, can help improve the 3D spatial reasoning abilities of LMMs. We consider two series of open-sourced LMMs, QWen2.5 [7] and InternVL2.5 [16], with a range of language model sizes from 0.5B to 72B. From the results shown in Fig. 4, we see that the scaling of language model sizes can effectively improve the 3D spatial reasoning abilities of LMMs. Larger language models with more parameters exhibit enhanced reasoning abilities. They can better capture 3D-aware information from the visual features and perform more complicated 3D spatial reasoning. However, given the importance of 3D spatial reasoning in a broad range of downstream applications, scaling up language model size is highly inefficient — LMMs with over 70B parameters exceed the computation capacity of common robotics or embodied AI systems and significantly limit the model throughput.
Figure 4: Scaling of language model sizes. Results show that scaling language model sizes can effectively improve the 3D spatial reasoning abilities of LMMs. However, with a 72B language model and a 6B vision encoder, InternVL2.5 still falls far behind human-level performance by more than 40%.
@@ 4.3 Robustness to Uncommon Camera Viewpoints
(a) Failure case of GPT-4o [29].
(b) Failure case of Gemini 2.0 Flash thinking [23].
Figure 5: Failures cases of GPT-4o [29] (top) and Gemini 2.0 Flash thinking [23] (bottom) on our 3DSRBench. (a) GPT-4o: GPT-4o does not have an explicit 3D representation, e.g., metric depth, and resort to visual cues to compare the distance, which leads to a wrong answer. (b) Gemini 2.0 Flash thinking: In this example Gemini 2.0 Flash thinking successfully breaks down the 3D spatial reasoning question into small and tractable steps. However, without explicit 3D representations, the model cannot perform reliable 3D spatial reasoning and predicts a wrong answer. See Sec. E in supplementary materials for more failure cases of the two models.
We study the robustness of 3D spatial reasoning abilities w.r.t. common and uncommon viewpoints. We evaluate a variety of open-sourced and proprietary LMMs on our 3DSRBench-synthetic-common and 3DSRBench-synthetic-uncommon splits and analyze the relative performance drop, given by
$\delta=\frac{\text{Acc}_{\text{uncommon}}-\text{Acc}_{\text{common}}}{\text{Acc}_{\text{common}}}$
As demonstrated by the results in Tab. 4, all LMMs exhibit significantly degraded performance when generalizing from common to uncommon viewpoints, e.g., a 13.5% drop in accuracy for GPT-4o [29], a 32.2% drop for Gemini-1.5-Pro [51], and a 19.1% drop for LLaVA-NeXT-8B [40]. We visualize two failure examples of GPT-4o in Fig. 3, showing how it is capable of predicting the correct answer when prompting with an image from a common viewpoint but fails when asked the same question with an image rendered from an uncommon viewpoint of the exact same scene.
We attribute such degraded performance in uncommon viewpoints to two factors: (i) image domain gap due to different camera viewpoint distributions between LMM training data and our 3DSRBench-synthetic-uncommon, and (ii) state-of-the-art LMMs adopt an implicit representation of 3D scenes. They are heavily built on the scaling law of data-driven approaches and lack explicit 3D representations that enable reliable 3D spatial reasoning. Despite the success of data-driven methods on a range of academic and empirical benchmarks, they face severe challenges generalizing to less represented data, which in our case, are images from uncommon camera 6D viewpoints.
These findings show that 3D spatial reasoning abilities of state-of-the-art LMMs are not robust to uncommon camera viewpoints. This largely limits their applicability to various downstream applications in robotics and embodied AI. Cameras mounted on robot arms or embodied AI systems are often positioned in uncommon locations and orientations as used in our study (see Fig. 3). On the one hand impressive advancements achieved by state-of-the-art LMMs in standard spatial reasoning benchmarks [30, 28, 36] may not generalize to downstream tasks; on the other hand, significantly degraded performance in uncommon viewpoints raises serious concerns about AI safety [3].
@@ 4.4 Failure Cases
We present two failure cases of GPT-4o [29] and Gemini 2.0 Flash thinking [23] in Fig. 5. In Fig. 5(a) we see that GPT-4o cannot perform rigorous 3D spatial reasoning and resort to various visual cues for reasoning. This is because GPT-4o lacks explicit 3D representations, e.g., metric depth, that limits its ability to perform complex 3D spatial reasoning. In Fig. 5(b), Gemini 2.0 Flash thinking successfully breaks down the 3D reasoning question into small and tractable steps. However, without explicit 3D representations, the model cannot perform reliable 3D spatial reasoning step-by-step. Despite the good thinking, the model fails to follow the planning and predicts a wrong answer.
We argue that for 3D spatial reasoning problems, models must not only have strong visual encoders to parse 3D-aware features, but also build a powerful reasoning model on various 3D information. Although scaling language model size leads to stronger reasoning abilities (see Fig. 4), a lack of explicit 3D representations would fundamentally limit models’ abilities to solve complex 3D spatial reasoning questions that require multi-step 3D computations.
@@ 5 Conclusions
In this work we study the 3D spatial reasoning capabilities of LMMs. We introduce a new benchmark, 3DSRBench, by manually annotating 2,100 visual question-answer pairs on natural images from MS-COCO, featuring diverse and open-vocabulary entities and a balanced data distribution for robust evaluation. To study the robustness of 3D spatial reasoning capabilities w.r.t. camera 6D viewpoints, we further annotate 672 visual question-answer pairs on synthetic multi-view images, each with a common and an uncommon camera viewpoint. We benchmark a wide variety of open-sourced and proprietary LMMs on our 3DSRBench, studying various 3D spatial reasoning capabilities, e.g., height, location, orientation, and multi-object reasoning, as well as the robustness of these LMMs to uncommon camera viewpoints. We also study how various designs of visual encoders and scaling of language models benefit 3D spatial reasoning. Experimental results on 3DSRBench provide valuable findings and insights to develop LMMs with strong 3D spatial reasoning abilities, as well as selecting LMMs for downstream applications that require robust 3D spatial reasoning.
B Baseline Models
@@
To analyze the 3D spatial reasoning capabilities of state-of-the-art LMMs, we explore a variety of proprietary LMMs, e.g., QWen-VL [7], Claude [4], Gemini [23], and GPT-4o [29]. For Gemini-2.0-Flash, we further evaluate two variants: (1) for Gemini-2.0-Flash-bbox we first ask the model to detect the 3D bounding box and then prompt the model to answer the question based on the 3D bounding boxes, and (2) for Gemini-2.0-Flash-think we evaluate the thinking model where the model performs deep thinking prior to answering the question.
@@
C Qualitative Examples of 3DSRBench
@@
D Qualitative Examples of Common and Uncommon Viewpoints
We present some qualitative examples of 3DSRBench-synthetic with multi-view images rendering the same scene and objects but from common (left) and uncommon (right) viewpoints in Fig. 9.
E Failure Cases
We present some failures cases of GPT-4o [29] in Fig. 10 and of Gemini 2.0 Flash thinking [23] in Fig. 11.
@@ E Failure Cases
Figure 9: Qualitative examples of multi-view images rendering the same scene and objects but from common (left) and uncommon (right) viewpoints.
Figure 10: Failure cases of GPT-4o [29] on our 3DSRBench. We find that GPT-4o cannot perform rigorous 3D spatial reasoning. In the top figure, GPT-4o resort to visual cues for 3D spatial reasoning. In the bottom two examples, GPT-4o understands the question, attempts to perform 3D spatial reasoning from the scene, and fails to derive the correct answer.
Figure 11: Failure cases of Gemini 2.0 Flash thinking [23] on our 3DSRBench. In this example Gemini 2.0 Flash thinking again breaks down the 3D spatial reasoning question into small and tractable steps. It also derives visual cues that can help verify the answer. However, the model still fails to perform 3D spatial reasoning over multiple 3D information, e.g., depth and orientation, and predicts a wrong answer.